Shift report 20000825 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Fri Aug 25 2000 - 00:06:29 EDT

  • Next message: http account: "Shift report 20000825 00:00-08:00"

    ------------------------------------------------------------
    Date: Fri Aug 25 00:06:28 2000, Shift: 16:00-24:00
    Supervisor: JJG, with: ZM, EJ
    Log book#5 pages 9 - 13
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam stored cogged at 17.15. Y=20000 and B=12000.
    2. Running nicely at high rate ZDC coinc. approx. 150/sek.
       Polarity: MRS at polarity A and FS at polarity B.
       Good beam since 6 pm and still running. All runs with
       D1 polarity B (625A), D2 polarity B (492A) and D5 with
       polarity A (350A). 
    
       The following numbers are measured with trigger 1. Runs
       2428 to 2431 are various short runs with magnets on/off
       Run 2428 has  17265 ( 26516 events)
       Run 2430 has 160336 (254999 events)
       Run 2431 has  33274 ( 52987 events)
       Run 2432 has 177642 (288680 events)
       Run 2433 has 216476 (362339 events)
       Run 2434 has 258188 (445297 events)
     
    3. T1 running nicely at Va=+1600; Vf=+700
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Fri Aug 25 2000 - 00:10:10 EDT